# Function Induction and Task Generalization: An Interpretability Study with Off-by-One Addition

### v1

@@ 4 Circuit Evaluation and Analysis
In Fig. 5, we present these heatmaps for each of the six FI heads identified in §3.2. We include three additional heads (H32.4, H28.6, H24.9) that, while showing a weaker effect (i.e., $1\%<|r|<2\%$) in §3.2, exhibit notable pattern in this analysis. We find that each of the FI heads contributes a distinct fragment to the overall +1 function. For example, with an input $x$, H39.7 promotes $x+1$, H28.6 suppresses $x-1$, H32.1 promotes digits greater than $x$, H24.9 suppress $x$. When the outputs of these nine heads are aggregated and added to the final residual stream, their combined effect implements the +1 function, as depicted in the bottom-right panel of Fig. 5.
Universality of Function Induction.
To investigate the universality of our findings across models, we repeat the path patching experiments with Llama-3 (8B) and Mistral-v0.1 (7B). We identified all three groups of heads in Llama-3 (8B) that account for their behavior in off-by-one addition. For Mistral-v0.1 (7B), we only identified FI heads and PT heads, suggesting a slight variation. Still, these observations provide promising evidence that the function induction mechanism is general and consistently emerges across various language models. See §C.2 for more details.
5 Task Generalization with Function Induction
Our investigation so far suggests that function induction is the key mechanism enabling the model to generalize from standard addition and manage the unexpected +1 step in off-by-one addition. Given that task generalization is crucial for developing capable AI systems, we aim to explore the broader usage of this mechanism. In this section, we investigate the role of function induction in a range of synthetic and algorithmic tasks. Specifically, §2 introduces the four task pairs examined, and §5.2 presents the overall findings and additional analyses for two of these pairs.
5.1 Tasks
(a) Off-by-$k$ Addition (c) Caeser Cipher
@@ 5.1 Tasks
(a) Off-by-$k$ Addition.
One extension of off-by-one addition is changing the offset to other values. Here, we consider offsets $k\in\{-2,-1,2\}$. We use $k=2$ as a representative case to be reported in the main paper. Results and analysis on the other offsets are deferred to §E.
(b) Shifted Multiple-choice QA.
@@
(c) Caesar Cipher.
One realistic task that leverages shifting functions is Caesar Cipher. During encoding, a letter is replaced by the corresponding letter a fixed number of positions down the alphabet [44]. This task is also commonly used to evaluate a language model’s reasoning capabilities [31]. Here we consider single-character Ceaser Cipher with different offsets $k\in\{-12,-11,\dots,0,\dots,12,13\}$. We use $k=2$ as the representative case in the main paper.
(d) Base-$k$ Addition.
Lastly, we consider the task of base-$k$ addition, which was used by Wu et al. [47] to assess the a model’s memorization versus generalization. Prior work [50] suggests that LMs may formulate a shortcut solution for base-8 addition by interpreting it as “adding 22 to the sum” from in-context examples; our interpretability analysis helps further investigate this observation. We consider two digit base-10 addition as the base task, and base-$k$ addition as the contrast task, with $k\in\{6,7,8,9\}$. We use $k=8$ as a representative case in the main paper.
@@ 5.2 Results and Analysis
FI heads are reused in a wider range of tasks.
Using the four task pairs introduced previously, we examine the role of the function induction mechanism we discover with head ablation experiments, similar to the one done in Fig. 4(a). We run forward passes on both the base task and the contrast task. We then replace the FI heads outputs in $M(.|x_{cont})$ forward pass with the corresponding head outputs in the $M(.|x_{base})$ forward pass.
We report results of the representative cases in Fig. 6. In all four task pairs, we first see a non-trivial performance on the contrast task, indicating effective generalization. Upon ablating the six FI heads, we observe a consistent trend: the model’s contrast accuracy significantly decreases; the base accuracy increases and often returns to a level comparable to that achieved with the base prompt. These findings suggest that the mechanism identified with off-by-one addition is largely reused in these task pairs, which share a similar underlying structure but also represents substantially different sub-steps. This strongly demonstrates the mechanism’s flexibility and composability.
We also observe that in (b) Shifted MMLU and (c) Caesar Cipher, the model has non-zero contrast accuracies when the FI heads are ablated. This implies that the six FI heads we found with off-by-one addition are useful, but not complete for these task pairs. See §E for additional discussion.
Figure 6: Task Generalization with FI Heads. (d) Base-8 addition has non-zero accuracy when using the base prompt, as some correct answers are identical in both base-10 and base-8.
Function vector analysis with off-by-$k$ addition.
We revisit the function vector style analysis done in Fig. 5, but this time considering different offsets $k\in\{-2,-1,1,2\}$. Results on two representative heads (H39.7 and H25.13) are shown in Fig. 7. We find that the effect of FI heads varies appropriately with the offset $k$, demonstrating their generality and consistency with the hypothesized functionality.
For these two selected heads, we find that each of them has their own “specialty.” For example, the heatmap for H25.13 suggests its primary responsibility for writing out ±$\pm 2$ functions. While its effect is stronger when the offset $k=\pm 2$, it still contributes in the case of $k=\pm 1$ by suppressing the original output $x$.
@@
Models struggle in base-8 addition due to under- or over-generalization.
It may sound unintuitive why the induction mechanism specialized in shifting functions could facilitate base-8 addition. One possible explanation is that the model initially performs standard base-10 addition with early layers, and apply minor adjustments when necessary. This adjustment step is possibly handled by the function induction mechanism in late layers.
@@
Case 1 93 6 1 0 100
Case 2 68 0 16 16 100
Case 3 83 14 0 0 100
Table 3: Error analysis for two-digit base-8 addition. We use 100 examples for each case specified in Listing 1. The anticipated behavior is marked in green. “Neither” suggests the number of times that neither c[0] or c[1] is adjusted, which is anticipated in Case 1. “c[0]” suggests that only c[0] is adjusted. “Both” suggests both digits are adjusted.
@@ 6 Related Works
Mechanistic Interpretability.
The field of mechanistic interpretability aims to reverse-engineer complex neural networks into human-understandable algorithms [2, 34], enhancing our understanding of a wide range of model behaviors, including in-context learning [30], long-context retrieval [46], and chain-of-thought reasoning [4]. A common methodology involves analyzing their computation graphs of a specific task, as exemplified by studies on indirect object identification [42], “greater than” operation [11], and entity tracking [32]. Following this, our work begins with the off-by-one addition task, and showcases the broader applicability of our findings with various task pairs.
Function Vectors in LMs.
Recent work has characterized in-context learning in language models as the compression of in-context examples into a single task or function vector, which is subsequently transported to the test example to trigger the model to apply the function [37, 12, 51]. These studies present strong evidence pertaining to single-step, mapping-style tasks like country-to-capital and English-French translation. Our work is inspired by this line of research, yet with two key differences: (1) We focus on off-by-one addition, a multi-step arithmetic task, where the learning of the second step depends on the results of the preceding step. (2) We provide a finer-grained interpretation on how function vectors, sent out by different attention heads, vary in content but collaborate to form a complete function. In concurrent work, this latter aspect was also explored by Hu et al. [14], who investigate the task of add-$k$ (i.e., “5$\rightarrow$8, 1$\rightarrow$4, 2$\rightarrow$?”) using subspace decomposition.
Latent Multi-step Reasoning and Structural Compositionality in LMs.
Various studies investigate whether and how models perform latent multi-step reasoning, typically via multi-hop factoid QA tasks [49, 41]. Our work demonstrates that LMs can dynamically infer the second step in a multi-step problem from in-context examples, a process representing a novel, flexible and composable form of latent multi-step reasoning. More broadly, our findings are relevant to research investigating structural compositionality [18] (i.e., breaking down complex tasks into subroutines) in language models.
@@ 7 Conclusion
In this work, we present an interpretability study on the off-by-one addition task, with the broader goal of investigating how language models handle unseen tasks using in-context learning. Our analysis led to the discovery of a function induction mechanism, which handles the “twists” involved in generalizing from seen to unseen tasks. This discovery extends and generalizes previous interpretability findings on induction heads and function vectors. We further show this mechanism is broadly reused beyond off-by-one addition, notably in realistic algorithmic tasks like Caesar Cipher and base-8 addition. Collectively, these observations deepen our understanding of how models generalize to new tasks and situations, and provide compelling evidence that language models may have developed composable and general mechanisms for handling intricate linguistic and task structures.
@@ Acknowledgment
Appendix A Induction Head Mechanism and Function Induction Mechanism
@@
Fig. 8 provides a side-by-side visualization of the induction head mechanism [30] and the hypothesized function induction mechanism (§3.2), demonstrating their structural similarity and explaining the basis for our hypothesis.
@@
Various prior works investigate how language models handle false, random, or perturbed demonstrations in in-context learning [25, 52, 43, 21, 20]. Notably, Halawi et al. [10] adopted an interpretability approach, observing the overthinking behavior of models (i.e., models draft truthful answers at early layers and flip them to untruthful answers at late layers), and identified false induction heads that are responsible for copying the untruthful answers from the ICL examples.
Our analysis of off-by-one addition was largely motivated by these studies. Here we revisit the findings of Halawi et al. [10] along with ours, using a unified view of two-step tasks, i.e., $z=f(g(x))$. In [10], the first step, $y=g(x)$ is typically a text classification task, e.g., news topic classification, and the second step, $z=f(y)$ is a permutation of the labels, e.g., {Business→Sci/Tech, Sci/Tech→World, World→Sport, Sports→Business}. In our work, $y=g(x)$ is standard addition, and $z=f(y)$ is a +1 function.
In this view, our findings with off-by-one addition are consistent with those in [10], while also advancing the understanding in several aspects: (1) In both cases, language models decompose the task into two steps, and induce the second step based on the results of the first step. The second step could be either a conditional copy-paste function, e.g., a permutation of labels, or an algorithmic function, e.g., a +1 function. The latter represents a novel finding of this study, demonstrating that the second step can exhibit forms more complex than copy-paste operations. (2) Our path patching procedure identified two additional group of heads ( consolidation heads and previous token heads) that are involved in handling false demonstrations. (3) Our work also suggests that the strategy to improve truthfulness by zeroing out false induction heads or function induction heads may have unintended consequences on models’ positive capabilities, given their positive contributions to the cipher task and the base-8 addition task.
Related to the view of two-step tasks, Jain et al. [15] demonstrate that models learn a “wrapper” function $g$ over an existing function $f$ in a sequential fine-tuning setting. Our work and [10] suggest that language models demonstrate simple forms of this behavior with in-context learning as well.
Appendix B Off-by-One Addition Evaluation
@@
Previously in Fig. 1, we reported results when the operands were sampled from the range of [0,999]. In Fig. 10, we additionally report results when sampling from the range of [0,9] and [0,99]. For two models using 0-9 tokenization (Gemma-2 (9B) and Qwen-2.5 (7B)), the performance drops with larger number ranges. For the remaining models, the performance remains stable regardless of the number ranges. [^4] We chose Gemma-2 (9B) as the default model in our study because (1) we focused on the range of [0,9] in early stage of this work to prioritize simplicity, and Gemma-2 (9B) performs competitively in this setting; (2) Qwen-2.5 (7B), Llama-3 (8B), Phi-4 (14B) were not released or integrated into transformer-lens at that time. We acknowledge this experimental design limitation and address it by interpreting Llama-3 (8B) and Mistral-v0.1 (7B) in §C.
@@
In Fig. 12-13, we present 32-shot off-by-$k$ addition results with various offsets $k$ using Gemma-2 (9B) and Llama-3 (8B) respectively. [^5] The visualization of Fig. 12-13 was inspired by Prabhakar et al. [31]. One consistent trend is that models struggle more with offsets $k$ of larger absolute values. While Llama-3 (8B) generally outperforms Gemma-2 (9B), Gemma-2 (9B) demonstrates strong performance when $k=\pm 10$, potentially due to its adoption of 0-9 tokenization. An additional observation reveals that Gemma-2 (9B) typically achieves stronger performance with even values of $k$ compared to odd values.
Appendix C Circuit Discovery
C.1 Relative Logit Diff
@@
C.2 Identified Heads
In the main paper, we focus on interpreting Gemma-2 (9B). To explore the universality of the mechanism, we additionally conduct path patching with Llama-3 (8B) and Mistral-v0.1 (7B). We list the identified attention heads below.
C.2.1 Gemma-2 (9B)
@@ C.2.2 Llama-3 (8B)
Figure 15: Circuit Discovery with Mistral-v0.1 (7B). Results are mostly consistent with those with Gemma-2 (9B) in Fig. 2, with the exception of the consolidation heads.
@@
C.2.3 Mistral-v0.1 (7B)
Mistral-v0.1 (7B) has 32 layers and 32 heads per layer. Path patching experiments were conducted with 4-shot off-by-one addition with numbers sampled from range [0,9]. We visualize the results in Fig. 15. For the two consolidation heads in the list below, they show weaker effect and attend to both the current token and some other tokens, which slightly deviates from our findings with Gemma-2 (9B) and Llama-3 (8B). Apart from this, the results using Mistral-v0.1 are consistent with our function induction hypothesis.
@@
C.3 Additional Interpretability Analysis
C.3.1 Logit Lens Analysis

### v2

@@ 5.1 Tasks
(a) Off-by-$k$ Addition. One extension of off-by-one addition is changing the offset to other values. Here, we consider offsets $k\in\{-2,-1,2\}$. We use $k=2$ as a representative case to be reported in the main paper. Results and analysis on the other offsets are deferred to §G.
@@
(c) Caesar Cipher. One realistic task that leverages shifting functions is Caesar Cipher. During encoding, a letter is replaced by the corresponding letter a fixed number of positions down the alphabet (Wikipedia contributors, 2025). This task is also commonly used to evaluate a language model’s reasoning capabilities (Prabhakar et al., 2024). Here we consider single-character Ceaser Cipher with different offsets $k\in\{-12,-11,\dots,0,\dots,12,13\}$. We use $k=0$ as the base task, and $k=2$ as the representative contrast task.
(d) Base-$k$ Addition. Lastly, we consider the task of base-$k$ addition, which was used by Wu et al. (2024) to assess the a model’s memorization versus generalization. Prior work (Ye et al., 2024) suggests that LMs may formulate a shortcut solution for base-8 addition by interpreting it as “adding 22 to the sum” from in-context examples; our interpretability analysis helps further investigate this observation. We consider two digit base-10 addition as the base task, and base-$k$ addition as the contrast task, with $k\in\{6,7,8,9\}$. We use $k=8$ as a representative case in the main paper.
@@ 5.2 Results and Analysis
FI heads are reused in a wider range of tasks. Using the four task pairs introduced above, we examine the role of the function induction mechanism we discover with head ablation experiments, similar to the one done in Fig. 4. We run forward passes on both the base task and the contrast task. We then replace the FI heads outputs in $M(.|x_{cont})$ forward pass with the corresponding head outputs in the $M(.|x_{base})$ forward pass.
We report results of the representative cases in Fig. 6. In all four task pairs, we first see a non-trivial performance on the contrast task, indicating effective generalization. Upon ablating the six FI heads, we observe a consistent trend: the model’s contrast accuracy substantially decreases; the base accuracy increases and often returns to a level comparable to that achieved with the base prompt. These findings suggest that the mechanism identified with off-by-one addition is largely reused in these task pairs, which share a similar underlying structure but also represents substantially different sub-steps. This strongly demonstrates the mechanism’s flexibility and composability.
We also observe that in (b) Shifted MMLU and (c) Caesar Cipher, the model has non-zero contrast accuracies when the FI heads are ablated. This implies that the six FI heads we found with off-by-one addition are useful, but not complete for these task pairs. See §G for additional discussion.
Figure 6: Task Generalization with FI Heads. In (d), base-8 addition has non-zero accuracy with the base-10 prompt, because in these test cases the base-10 answers happen to be correct in base-8.
Function vector analysis with off-by-$k$ addition. We revisit the function vector style analysis done in Fig. 5, but this time considering different offsets $k\in\{-2,-1,1,2\}$. Results on two representative heads (H39.7 and H25.13) are shown in Fig. 7, with other heads deferred to Fig. 23-25.
We find that the effect of FI heads varies meaningfully with the offset $k$, demonstrating their generality and consistency with the hypothesized functionality. For the two selected heads in Fig. 7, we find that each of them has their own “specialty.” For example, the heatmap for H25.13 suggests its primary responsibility for writing out $\pm 2$ functions. While its effect is stronger when the offset $k=\pm 2$, it still contributes in the case of $k=\pm 1$ by suppressing the original output $x$.
@@
Models struggle in base-8 addition due to under- or over-generalization. It may sound unintuitive why the induction mechanism specialized in shifting functions could facilitate base-8 addition. One possible explanation is that the model initially performs standard base-10 addition with early layers, and apply minor adjustments when necessary. This adjustment step is possibly handled by the function induction mechanism in late layers.
@@
[…]
Table 3: Error analysis for two-digit base-8 addition. We use 100 examples for each case specified in Listing 1. The correct behavior is marked in green. “Neither” suggests the number of times that neither c[0] or c[1] is adjusted, which is anticipated in Case 1. “c[0]” suggests that only c[0] is adjusted. “Both” suggests both digits are adjusted.
@@ 6 Related Works
Mechanistic Interpretability. The field of mechanistic interpretability aims to reverse-engineer complex neural networks into human-understandable algorithms (Bereska & Gavves, 2024; Sharkey et al., 2025), enhancing our understanding of a wide range of model behaviors, including in-context learning (Olsson et al., 2022), long-context retrieval (Wu et al., 2025), and chain-of-thought reasoning (Cabannes et al., 2024). A common methodology involves analyzing their computation graphs of a specific task, as exemplified by studies on indirect object identification (Wang et al., 2023), “greater than” operation (Hanna et al., 2023), and entity tracking (Prakash et al., 2024a). Following this, our work begins with the off-by-one addition task, and showcases the broader applicability of our findings with various task pairs.
[…]
Latent Multi-step Reasoning and Structural Compositionality in LMs. Various studies investigate whether and how models perform latent multi-step reasoning, typically via multi-hop factoid QA tasks (Yang et al., 2024b; Wang et al., 2024). Our work demonstrates that LMs can dynamically infer the second step in a two-step problem from in-context examples, a process representing a novel, flexible and composable form of latent multi-step reasoning. More broadly, our findings are relevant to research investigating structural compositionality (Lepori et al., 2023) (i.e., breaking down complex tasks into subroutines) in language models.
@@ 7 Conclusion
In this work, we present an interpretability study on the off-by-one addition task, with the broader goal of investigating how language models handle unseen tasks using in-context learning. Our analysis led to the discovery of a function induction mechanism, which handles the “twists” involved in generalizing from seen to unseen tasks. This discovery extends and generalizes previous interpretability findings on induction heads and function vectors. We further show this mechanism is broadly reused beyond off-by-one addition, notably in realistic algorithmic tasks like Caesar Cipher and base-8 addition. Collectively, these observations deepen our understanding of what language models are capable of with in-context learning and multi-step reasoning, and how models generalize to novel tasks and situations. Moreover, our work provides compelling evidence that language models may develop composable and general mechanisms for handling ever-changing task variations, suggesting one possible pathway toward explaining and perhaps further enhancing model capabilities.
@@ Acknowledgment
Appendix A Limitations
Regarding circuit discovery experiments (§3, §D and §E), the identified circuit is limited as it does not perfectly satisfy the faithfulness and completeness criteria, even with our best efforts. This challenge arises because achieving simultaneous satisfaction of faithfulness, completeness, and minimality is difficult, as these criteria often regulate each other. Moreover, number tokens are often mapped into a sinusoidal (Fourier) feature space rather than a linear space in language models (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024), which further complicates our interpretability analysis.
Regarding circuit analysis (§4 and §F), we mainly used causal intervention methods and examine the causal effect of attention heads on naive prompts. Future work could provide deeper mechanistic insights by analyzing the query-key and output-value circuits within these heads (Elhage et al., 2021), or by investigating the role of MLP layers in the overall mechanism.
Regarding task generalization experiments (§5 and §G), our current scope is limited to two-step tasks where the second step involves a shifting-related function. We anticipate that the function induction mechanism could operate on a broader spectrum of functions, which could be investigated in future work. Additionally, the task pairs we investigated are synthetic or algorithmic; further exploration of the role of function induction heads on naturally occurring texts would be highly valuable.
Appendix B Induction Head Mechanism and Function Induction Mechanism
@@
Fig. 8 provides a side-by-side visualization of the induction head mechanism (Olsson et al., 2022) and the hypothesized function induction mechanism (§3.2), demonstrating their structural similarity and explaining the basis for our hypothesis.
@@
Various prior works investigate how language models handle false, random, or perturbed demonstrations in in-context learning (Min et al., 2022b; Yoo et al., 2022; Wei et al., 2024; Lyu et al., 2023; Lin & Lee, 2024). Notably, Halawi et al. (2023) adopted an interpretability approach, observing the overthinking behavior of models (i.e., models draft truthful answers at early layers and flip them to untruthful answers at late layers), and identified false induction heads that are responsible for copying the untruthful answers from the ICL examples.
Our analysis of off-by-one addition was largely motivated by these studies. Here we revisit the findings of Halawi et al. (2023) along with ours, using a unified view of two-step tasks, i.e., $z=f(g(x))$. In Halawi et al. (2023), the first step, $y=g(x)$ is typically a text classification task, e.g., news topic classification, and the second step, $z=f(y)$ is a permutation of the labels, e.g., {Business→Sci/Tech, Sci/Tech→World, World→Sport, Sports→Business}. In our work, $y=g(x)$ is standard addition, and $z=f(y)$ is a +1 function.
In this view, our findings with off-by-one addition are consistent with those in Halawi et al. (2023), while also advancing the understanding in several aspects: (1) In both cases, language models decompose the task into two steps, and induce the second step based on the results of the first step. The second step could be either a conditional copy-paste function, e.g., a permutation of labels, or an algorithmic function, e.g., a +1 function. The latter represents a novel finding of this study, demonstrating that the second step can exhibit forms more complex than copy-paste operations. (2) Our path patching procedure has led us to identify two additional group of heads ( consolidation heads and previous token heads) that are involved in handling false demonstrations. (3) Our work also suggests that the strategy to improve truthfulness by zeroing out false induction heads or function induction heads may have unintended consequences on models’ positive capabilities, given their positive contributions to the cipher task and the base-8 addition task.
Related to the view of two-step tasks, Jain et al. (2024) demonstrate that models learn a “wrapper” function $g$ over an existing function $f$ in a sequential fine-tuning setting. Our work and Halawi et al. (2023) suggest that language models demonstrate simple forms of this behavior with in-context learning as well.
Appendix C Off-by-One Addition Evaluation
@@
Previously in Fig. 1, we reported results when the operands were sampled from the range of [0,999]. In Fig. 10, we additionally report results when sampling from the range of [0,9] and [0,99]. For two models using 0-9 tokenization (Gemma-2 (9B) and Qwen-2.5 (7B)), the performance drops with larger number ranges. For the remaining models, the performance remains stable regardless of the number ranges. [^4] We chose Gemma-2 (9B) as the default model in our study because (1) we focused on the range of [0,9] in early stage of this work to prioritize simplicity, and Gemma-2 (9B) performs competitively in this setting; (2) Qwen-2.5 (7B), Llama-3 (8B), Phi-4 (14B) were not released or integrated into transformer-lens at that time. We acknowledge this experimental design limitation and address it by interpreting Llama-3 (8B) and Mistral-v0.1 (7B) in §D.
@@
In Fig. 12-13, we present 32-shot off-by-$k$ addition results with various offsets $k$ using Gemma-2 (9B) and Llama-3 (8B) respectively. [^5] The visualization of Fig. 12-13 was inspired by Prabhakar et al. (2024). One consistent trend is that models struggle more with offsets $k$ of larger absolute values. While Llama-3 (8B) generally outperforms Gemma-2 (9B), Gemma-2 (9B) demonstrates strong performance when $k=\pm 10$, potentially due to its adoption of 0-9 tokenization. An additional observation reveals that Gemma-2 (9B) typically achieves stronger performance with even values of $k$ compared to odd values.
Appendix D Circuit Discovery
D.1 Relative Logit Diff
@@
D.2 Identified Heads
In the main paper, we focus on interpreting Gemma-2 (9B). To explore the universality of the mechanism, we additionally conduct path patching with Llama-3 (8B), Llama-2 (7B) and Mistral-v0.1 (7B). We list the identified attention heads below.
D.2.1 Gemma-2 (9B)
Gemma-2 (9B) has 42 layers and 16 heads per layer. Path patching experiments were conducted with 4-shot off-by-one addition with numbers sampled from range [0,9].
@@ D.2.2 Llama-3 (8B)
D.2.3 Mistral-v0.1 (7B)
Figure 15: Circuit Discovery with Mistral-v0.1 (7B). Results are mostly consistent with those with Gemma-2 (9B) in Fig. 2, with the exception of the consolidation heads showing weaker signals.
Mistral-v0.1 (7B) has 32 layers and 32 heads per layer. Path patching experiments were conducted with 4-shot off-by-one addition with numbers sampled from range [0,9]. We visualize the results in Fig. 15. For the two consolidation heads in the list below, they show weaker effect and attend to both the current token and some other tokens, which slightly deviates from our findings with Gemma-2 (9B). Apart from this, the results using Mistral-v0.1 are consistent with other models.
@@
D.2.4 Llama-2 (7B)
Figure 16: Circuit Discovery with Llama-2 (7B). Results are mostly consistent with those with Gemma-2 (9B) in Fig. 2.
Llama-2 (7B) has 32 layers and 32 heads per layer. Path patching experiments were conducted with 4-shot off-by-one addition with numbers sampled from range [0,9]. We visualize the results in Fig. 16.
All three groups of heads are present in Llama-2 (7B). However, we notice two small variations compared to the circuit in Gemma-2 (9B). (1) We identified H16.24 that achieves $r=2.12\%$, but its attention pattern doesn’t fit that of a consolidation head or a function induction head. Since $2.12\%$ is slightly above the $2\%$ threshold we set, we consider this noise; (2) One previous token head (H29.1) no longer attends to the “=” immediately before the answer token $c_{i}$ at the token $c_{i}$. Instead, it attends to the “=” token one ICL example away.
To discuss our findings with those of Todd et al. (2024) with a common ground, we extract the list of FV heads by selecting the 10 heads with the highest absolute average indirect effect (AIE) from Fig. 19 in Todd et al. (2024). These heads are concentrated in early-middle layers (before layer 20), whereas our FI heads appear in late layers (layers 29-31). There is no overlap between the two sets.
  • Consolidation Heads: H31.28, H31.10, H30.3;
  • Function Induction (FI) Heads: H31.30, H31.4, H29.26, H29.16, H30.26, H30.3;
  • Previous Token (PT) Heads: H30.13, H29.1, H28.5, H28.10, H28.16, H28.24, H27.31;
  • Miscellaneous Heads: H16.24;
  • Function Vector Heads (Todd et al., 2024): H9.25, H11.2, H11.18, H12.15, H12.18, H12.28, H13.7, H14.1, H14.16, H16.10.
D.3 Additional Interpretability Analysis
D.3.1 Logit Lens Analysis